1.The voltage on a transmission line with characteristic impedance 50 ΩΩis sin(31.4?+6.28×109?)sin⁡(31.4z+6.28×109t)V. What is the current associated with this wave?

A. cos(31.4?+6.28×109?)cos⁡(31.4z+6.28×109t)A

B. 0.02sin(31.4?+6.28×109?)0.02sin⁡(31.4z+6.28×109t)A

C. −.02sin(31.4?+6.28×109?)−.02sin⁡(31.4z+6.28×109t)A

D. 50sin(31.4?+6.28×109?)50sin⁡(31.4z+6.28×109t)A

E. None of the above

2.

The unit circle on a Smith chart represents:

Purely reactive impedances

Purely resistive impedances

Matched impedances

VSWR = 0

None of the above
Answer #1

1. B) 0.02sin(31.4?+6.28×109?)

I = V/Z = sin(31.4?+6.28×109?)/50 = 0.02sin(31.4?+6.28×109?)

2. None of the above.

The outter ring of the Smith Chart is where the magnitude of reflection coefficient is equal to 1, i.e. all energy is reflected. For matched impedance, there is no reflection i.e. = 0 and VSWR = 0. For || = 1, VSWR = . The unit circle does not represent purely resistive or reactive impedance. It only represents the impedance that causes maximum reflection w.r.t characteristic impedance.